OpenStudy (anonymous):

please help!!!! How do the graphs of y =1/x and y = 3/x – 4 compare? Thanks!

hero (hero):

Hint: Multiply 1/x by 3, then subtract 4.

OpenStudy (anonymous):

Compared to the graph of y =1/x , the graph of y =3/x – 4 is a vertical stretch by a factor of 3 and a translation of 4 units left. Compared to the graph of y =1/x , the graph of y = 3/x – 4 is a vertical shrink by a factor of 3 and a translation of 4 units up. Compared to the graph of y =1/x , the graph of y =3/x – 4 is a vertical stretch by a factor of 3 and a translation of 4 units down. d. Compared to the graph of y =1/x , the graph of y =3/x – 4 is a vertical shrink by a factor of 3 and a translation of 4 units down.

hero (hero):

You describe the graphs in this manner. You say: y = 3/x - 4 is the graph of 1/x amplified by 3 and shifted down by 4 units.

OpenStudy (anonymous):

so it would be the third one? :/

hero (hero):

Yup

OpenStudy (anonymous):

thank you!! <3